# A Parameter-Free and Near-Optimal Zeroth-Order Algorithm for Stochastic Convex Optimization

### v1

@@ 1 Introduction
Yes $\tilde{\mathcal{O}}\left(\dfrac{dL^{2}D_{\mathcal{X}}^{2}}{\epsilon^{2}}\right)$ $\dfrac{\bar{r}_{t}}{\sqrt{G_{t}}}$ $\sqrt{\dfrac{d}{t+1}}$
@@ 3 Parameter-Free Stochastic Zeroth-Order Optimization
4:$\mu_{t}=\sqrt{\dfrac{d}{t+1}}$
5:${\bf{v}}_{t}\sim{\mathbb{U}}({\mathbb{S}}^{d-1})$,   ${\bm{\xi}}_{t}\sim\Xi$
6:${\bf{g}}_{t}=\dfrac{d}{2\mu_{t}}(F({\bf{x}}_{t}+\mu_{t}{\bf{v}}_{t};{\bm{\xi}}_{t})-F({\bf{x}}_{t}-\mu_{t}{\bf{v}}_{t};{\bm{\xi}}_{t})){\bf{v}}_{t}$
7:$\bar{r}_{t}=\max\{\bar{r}_{t-1},\|{\bf{x}}_{t}-{\bf{x}}_{0}\|\}$
@@
$\mu_{t}\triangleq\sqrt{\frac{d}{t+1}},$ (6)
which is adaptive and larger than the smoothing parameters required in the convergence guarantees of existing stochastic zeroth-order methods [40, 49, 15, 19, 57, 45], e.g., Nesterov and Spokoiny [40], Shamir [49] takes $\mu_{t}={\mathcal{O}}(\sqrt{d/T}\,)$ in their analysis, which depends on the total iteration budget $T$; Duchi et al. [15] takes $\mu_{t}={\mathcal{O}}(1/(dt)^{2})$, which may be quite small for high dimensional problems. Recall the smoothing parameter $\mu_{t}$ appears in the denominator of the stochastic finite difference (2). Therefore, the larger $\mu_{t}$ makes the iteration more stable.
@@ 4 The Complexity Analysis
We show that POEM (Algorithm 1) achieves the near-optimal SZO complexity. The detailed proofs for results in this section are differed to Appendix B.
@@
The settings of soothing parameter (6) implies that $\mu_{k}={\mathcal{O}}(\sqrt{d/k}\,)$, then we can use an integral to establish an upper bound for the series to prove Lemma 6. We present the detailed proof in Appendix B.3.
@@
$f(\bar{{\bf{x}}}_{t})\!-\!f({\bf{x}}_{\star})\!\leq\!\frac{16(\theta_{t,\delta}(\bar{r}_{t}\!+\!s_{0})(\sqrt{G_{t-1}}\!+\!Ld)\!+\!L\sqrt{dt}\,)}{\sum_{k=0}^{t-1}\bar{r}_{k}/\bar{r}_{t}},$
@@ 5 Results for Unbounded Domains
with probability at least $1-\delta$, where $s_{0}=\|{\bf{x}}_{0}-{\bf{x}}_{\star}\|$ and $\alpha_{T,\delta}\triangleq\log_{+}(T+1)\log(60\log(T/\delta))$.
@@ B.4 Proof of Proposition 1
$\frac{(2\bar{s}_{t}+\bar{r}_{t})\sqrt{G_{t-1}}+8\bar{s}_{t}\sqrt{\theta_{t,\delta}G_{t-1}+4L^{2}d^{2}\theta_{t,\delta}^{2}}\,+4L\sqrt{dt}}{\sum_{k=0}^{t-1}\bar{r}_{k}/\bar{r}_{t}},$
@@
$\frac{(2\bar{s}_{t}+\bar{r}_{t})\sqrt{G_{t-1}}+8\bar{s}_{t}(\theta_{t,\delta}\sqrt{G_{t-1}}+2\theta_{t,\delta}Ld)+4L\sqrt{dt}}{\sum_{k=0}^{t-1}\bar{r}_{k}/\bar{r}_{t}}.$
@@
$16\cdot\frac{\theta_{t,\delta}(\bar{r}_{t}+s_{0})(\sqrt{G_{t-1}}+Ld)+L\sqrt{dt}}{\sum_{k=0}^{t-1}\bar{r}_{k}/\bar{r}_{t}}.$
@@ B.5 Proof of Theorem 1
$f(\bar{{\bf{x}}}_{t})-f({\bf{x}}_{\star})\leq 16\cdot\frac{\theta_{t,\delta}(\bar{r}_{t}+s_{0})(\sqrt{G_{t-1}}+Ld)+L\sqrt{dt}}{\sum_{k=0}^{t-1}\bar{r}_{k}/\bar{r}_{t}},$
@@
$f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})\leq c_{0}\cdot\frac{\theta_{\tau_{T},\delta}(\bar{r}_{\tau_{T}}+s_{0})(\sqrt{G_{\tau_{T}-1}}+Ld)+L\sqrt{dt}}{T}\log_{+}\bigg{(}\frac{\bar{r}_{\tau_{T}}}{r_{\epsilon}}\bigg{)},$
@@
$f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})\leq c_{0}\cdot\frac{\theta_{T,\delta}(\bar{r}_{T}+s_{0})(\sqrt{G_{T-1}}+Ld)+L\sqrt{dT}}{T}\log_{+}\bigg{(}\frac{\bar{r}_{T}}{r_{\epsilon}}\bigg{)}.$
@@
$f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})\leq c_{0}\cdot\frac{\theta_{T,\delta}D_{\mathcal{X}}(\sqrt{G_{T-1}}+Ld)+L\sqrt{dT}}{T}\log_{+}\bigg{(}\frac{D_{\mathcal{X}}}{r_{\epsilon}}\bigg{)}.$
@@
${\mathbb{E}}[f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})\,|\,{\mathcal{F}}_{\delta}]\leq c_{0}\cdot\frac{\theta_{T,\delta}D_{\mathcal{X}}({\mathbb{E}}[\sqrt{G_{T-1}}\,|\,{\mathcal{F}}_{\delta}]+Ld)+L\sqrt{dT}}{T}\log_{+}\bigg{(}\frac{D_{\mathcal{X}}}{r_{\epsilon}}\bigg{)}.$ (18)
@@ C.3 Proof of Theorem 2
We first establish the existence and uniqueness of the conditional expectation ${\mathbb{E}}[f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}})\,|\,\tilde{{\mathcal{F}}}_{\delta}]$. According to Durrett [16, Chapter 4.1], it suffices to verify that ~$\tilde{{\mathcal{F}}}_{\delta}\subset\tilde{{\mathcal{F}}}_{0}$ and ${\mathbb{E}}|f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})|<\infty$. By definition, we recall that $\tilde{F}_{\delta}=\{A:A\subset\tilde{\Omega}_{\delta}\,\cap\,\hat{\Omega}_{\delta}\}\cap\tilde{{\mathcal{F}}}_{0}$. Thus, we have ~$\tilde{{\mathcal{F}}}_{\delta}\subset\tilde{{\mathcal{F}}}_{0}$. Next, from inequality (14), we have

### v2

@@ 1 Introduction
Yes $\tilde{\mathcal{O}}\left(\dfrac{dL^{2}D_{\mathcal{X}}^{2}}{\epsilon^{2}}\right)$ $\dfrac{\bar{r}_{t}}{\sqrt{G_{t}}}$ $\bar{r}_{t}\sqrt{\dfrac{d}{t+1}}$
@@ 3 Parameter-Free Stochastic Zeroth-Order Optimization
4:$\bar{r}_{t}=\max\{\bar{r}_{t-1},\|{\bf{x}}_{t}-{\bf{x}}_{0}\|\}$
5:$\mu_{t}=\bar{r}_{t}\sqrt{\dfrac{d}{t+1}}$
6:${\bf{v}}_{t}\sim{\mathbb{U}}({\mathbb{S}}^{d-1})$,   ${\bm{\xi}}_{t}\sim\Xi$
7:${\bf{g}}_{t}=\dfrac{d}{2\mu_{t}}(F({\bf{x}}_{t}+\mu_{t}{\bf{v}}_{t};{\bm{\xi}}_{t})-F({\bf{x}}_{t}-\mu_{t}{\bf{v}}_{t};{\bm{\xi}}_{t})){\bf{v}}_{t}$
@@
$\mu_{t}\triangleq\bar{r}_{t}\sqrt{\frac{d}{t+1}},$ (6)
which is adaptive and larger than the smoothing parameters required in the existing stochastic zeroth-order methods [40, 49, 15, 19, 57, 45], e.g., Nesterov and Spokoiny [40], Shamir [49] takes $\mu_{t}={\mathcal{O}}(\sqrt{d/T}\,)$ in their analysis, which depends on the total iteration budget $T$; Duchi et al. [15] takes $\mu_{t}={\mathcal{O}}(1/(dt)^{2})$, which may be quite small for high dimensional problems. Recall the smoothing parameter $\mu_{t}$ appears in the denominator of the stochastic finite difference (2). Therefore, the larger $\mu_{t}$ makes the iteration more stable.
@@ 4 The Complexity Analysis
We show that POEM (Algorithm 1) achieves the near-optimal SZO complexity. The detailed proofs for results in this section are deferred to Appendix B.
@@
The settings of smoothing parameter (6) implies that $\mu_{k}={\mathcal{O}}(\sqrt{d/k}\,)$, then we can use an integral to establish an upper bound for the series to prove Lemma 6. We present the detailed proof in Appendix B.3.
@@
$f(\bar{{\bf{x}}}_{t})\!-\!f({\bf{x}}_{\star})\!\leq\!\frac{16\theta_{t,\delta}(\bar{r}_{t}\!+\!s_{0})(\sqrt{G_{t-1}}\!+\!Ld+\!L\sqrt{dt}\,)}{\sum_{k=0}^{t-1}\bar{r}_{k}/\bar{r}_{t}},$
@@ 5 Results for Unbounded Domains
with probability at least $1-2\delta$, where $s_{0}=\|{\bf{x}}_{0}-{\bf{x}}_{\star}\|$ and $\alpha_{T,\delta}\triangleq\log_{+}(T+1)\log(60\log(T/\delta))$.
@@ B.4 Proof of Proposition 1
$\frac{(2\bar{s}_{t}+\bar{r}_{t})\sqrt{G_{t-1}}+8\bar{s}_{t}\sqrt{\theta_{t,\delta}G_{t-1}+4L^{2}d^{2}\theta_{t,\delta}^{2}}\,+4\bar{r}_{t}L\sqrt{dt}}{\sum_{k=0}^{t-1}\bar{r}_{k}/\bar{r}_{t}},$
@@
$\frac{(2\bar{s}_{t}+\bar{r}_{t})\sqrt{G_{t-1}}+8\bar{s}_{t}(\theta_{t,\delta}\sqrt{G_{t-1}}+2\theta_{t,\delta}Ld)+4\bar{r}_{t}L\sqrt{dt}}{\sum_{k=0}^{t-1}\bar{r}_{k}/\bar{r}_{t}}.$
@@
$16\cdot\frac{\theta_{t,\delta}(\bar{r}_{t}+s_{0})(\sqrt{G_{t-1}}+Ld+L\sqrt{dt}\,)}{\sum_{k=0}^{t-1}\bar{r}_{k}/\bar{r}_{t}}.$
@@ B.5 Proof of Theorem 1
$f(\bar{{\bf{x}}}_{t})-f({\bf{x}}_{\star})\leq 16\cdot\frac{\theta_{t,\delta}(\bar{r}_{t}+s_{0})(\sqrt{G_{t-1}}+Ld+L\sqrt{dt}\,)}{\sum_{k=0}^{t-1}\bar{r}_{k}/\bar{r}_{t}},$
@@
$f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})\leq c_{0}\cdot\frac{\theta_{\tau_{T},\delta}(\bar{r}_{\tau_{T}}+s_{0})(\sqrt{G_{\tau_{T}-1}}+Ld+L\sqrt{dt}\,)}{T}\log_{+}\bigg{(}\frac{\bar{r}_{\tau_{T}}}{r_{\epsilon}}\bigg{)},$
@@
$f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})\leq c_{0}\cdot\frac{\theta_{T,\delta}(\bar{r}_{T}+s_{0})(\sqrt{G_{T-1}}+Ld+L\sqrt{dT}\,)}{T}\log_{+}\bigg{(}\frac{\bar{r}_{T}}{r_{\epsilon}}\bigg{)}.$
@@
$f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})\leq c_{0}\cdot\frac{\theta_{T,\delta}D_{\mathcal{X}}(\sqrt{G_{T-1}}+Ld+L\sqrt{dT}\,)}{T}\log_{+}\bigg{(}\frac{D_{\mathcal{X}}}{r_{\epsilon}}\bigg{)}.$
@@
${\mathbb{E}}[f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})\,|\,{\mathcal{F}}_{\delta}]\leq c_{0}\cdot\frac{\theta_{T,\delta}D_{\mathcal{X}}({\mathbb{E}}[\sqrt{G_{T-1}}\,|\,{\mathcal{F}}_{\delta}]+Ld+L\sqrt{dT}\,)}{T}\log_{+}\bigg{(}\frac{D_{\mathcal{X}}}{r_{\epsilon}}\bigg{)}.$ (18)
@@ C.3 Proof of Theorem 2
We first establish the existence and uniqueness of the conditional expectation ${\mathbb{E}}[f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})\,|\,\tilde{{\mathcal{F}}}_{\delta}]$. According to Durrett [16, Chapter 4.1], it suffices to verify that ~$\tilde{{\mathcal{F}}}_{\delta}\subset\tilde{{\mathcal{F}}}_{0}$ and ${\mathbb{E}}|f(\bar{{\bf{x}}}_{\tau_{T}})-f({\bf{x}}_{\star})|<\infty$. By definition, we recall that $\tilde{F}_{\delta}=\{A:A\subset\tilde{\Omega}_{\delta}\,\cap\,\hat{\Omega}_{\delta}\}\cap\tilde{{\mathcal{F}}}_{0}$. Thus, we have ~$\tilde{{\mathcal{F}}}_{\delta}\subset\tilde{{\mathcal{F}}}_{0}$. Next, from inequality (14), we have